OpenStudy (anonymous):

Medal to whoever can provide correct answer or help me If $240 is invested at an interest rate of 9% per year and is compounded monthly, how much will the investment be worth in 14 years? $68.39 $704.28 $842.13 $846.10

OpenStudy (anonymous):

A=P(1+rt)

OpenStudy (anonymous):

^that equation is going to help you a lot with these types of problems I can explain it in full, give me a second :)

OpenStudy (anonymous):

P is the principal (or starting amount) r is the interest rate, and t is the number of times it is compounded annually. So let's plug in the numbers

OpenStudy (mertsj):

\[A = P(1+\frac{r}{n})^{nt}\]

OpenStudy (mertsj):

\[A= 240(1+\frac{.09}{12})^{12(9)}\]

OpenStudy (anonymous):

^that is correct, sorry forgot that detail! but still we can just plug in our numbers for the answer.

OpenStudy (mertsj):

\[A=240(1+.0075)^{108}\]

OpenStudy (anonymous):

so your answer is 842.1325...

OpenStudy (anonymous):

which looks like C in this situation :) cheers!

OpenStudy (anonymous):

Thank you so much. Could yu help me with a few more questions like this? If it's too much trouble then it's okay :)

OpenStudy (anonymous):

Just keep in mind the formula A=P(1+r/n)^(n*t) P=principal r=interest rate n= number of times per year it is compounded (12 if monthly 4 if quarterly 2 if semiannual etc) t=number of years and it becomes quite simple, but yes I can help :)

OpenStudy (anonymous):

If $396 is invested at an interest rate of 13% per year and is compounded continuously, how much will the investment be worth in 3 years? $584.88 $583.66 $581.27 $268.11

OpenStudy (anonymous):

sorry, I'm at work, had a customer, but I'm back! Alright, so we can plug these numbers in for the answer.

OpenStudy (anonymous):

ooh, compounded continuously, so we may need to make a minor change to our equation. Usually when we see that something is compounded continuously we will use e (there is a button for e on your graphing calculator) A=Pe^rt but the rest is the same P = principal amount (initial investment) r = annual interest rate (as a decimal) t = number of years A = amount after time t

OpenStudy (anonymous):

so plug in to your graphing calc 396e^(.13*3)

OpenStudy (anonymous):

and you get 584.8843944, which means your answer would be A) $584.88